OpenStudy (celinegirl):

An expression is shown below: square root of 18 plus square root of 2 Which statement is true about the expression? A.) It is rational and equal to 3. B.)It is rational and equal to 4. C.)It is irrational and equal to 3 multiplied by the square root of 2. D.)It is irrational and equal to 4 multiplied by the square root of 2.

OpenStudy (malcolmmcswain):

Do you know how to set up this problem?

OpenStudy (malcolmmcswain):

@celinegirl

OpenStudy (malcolmmcswain):

Well, the 2 is already prime, so we don't have to factor it.

OpenStudy (malcolmmcswain):

We have to factor 18

OpenStudy (michele_laino):

hint: we can write this: \[\huge \begin{gathered} \sqrt {18} + \sqrt 2 = \sqrt {9 \cdot 2} + \sqrt 2 = \hfill \\ \hfill \\ = \sqrt 9 \cdot \sqrt 2 + \sqrt 2 = 3\sqrt 2 + \sqrt 2 = ...? \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

since \( \sqrt 9 = 3\)

OpenStudy (malcolmmcswain):

3 times the square root of two + 1 times the square root of two = ? @celinegirl

OpenStudy (malcolmmcswain):

And is that rational or irrational?

OpenStudy (celinegirl):

eh i dont know @malcolmmcswain